### v2

Abstract
We call a matrix blocky if, up to row and column permutations, it can be obtained from an identity matrix by repeatedly applying one of the following operations: duplicating a row, duplicating a column, or adding a zero row or column. Blocky matrices are precisely the boolean matrices that are contractive when considered as Schur multipliers. It is conjectured that any boolean matrix with Schur multiplier norm at most $\gamma$ is expressible as a signed sum
$A=\sum_{i=1}^{L}\pm B_{i}$
for some blocky matrices $B_{i}$, where $L$ depends only on $\gamma$. This conjecture is an analogue of Green and Sanders’s quantitative version of Cohen’s idempotent theorem. In this paper, we prove bounds on $L$ that are polylogarithmic in the dimension of $A$. Concretely, if $A$ is an $n\times n$ matrix, we show that one may take $L=2^{O(\gamma^{7})}\log(n)^{2}$. Keywords. Block complexity, Schur multipliers, factorization norm. MSC2020 Classification. 15B36, 47L80, 94D10.
1 Introduction
A celebrated classical result in harmonic analysis is the 1960 theorem of P. J. Cohen [Coh60], which states that a measure $\mu$ on a locally compact abelian group $G$ is idempotent if and only if $\widehat{\mu}$ is a member of the coset ring of $\widehat{G}$. This was conjectured by W. Rudin [Rud59], who settled the case of discrete groups and finite-dimensional torus groups. An explicit description of idempotent measures on the circle group was given earlier by H. Helson [Hel54]. Cohen’s idempotent theorem is sometimes referred to as the Cohen–Host idempotent theorem, as in 1986, B. Host extended Cohen’s result to arbitrary locally compact groups [Hos86].
For finite groups, Cohen’s theorem holds without offering any content, since in this case, the coset ring consists of all functions on $\widehat{G}$. However, a 2008 theorem of B. Green and T. Sanders [GS08b] gives a quantitative version of the idempotent theorem that is useful even for finite groups. Subsequent refinements by Sanders [San11, San20] give better bounds and show that the group need not even be abelian for this result to hold. For the sake of simplicity, in the brief exposition below, we shall take $G$ to be abelian and finite.
Let $\left|\!\left|f\right|\!\right|_{\rm A}$ denote the Fourier algebra norm $\sum_{a\in\widehat{G}}\bigl{|}\widehat{f}(a)\bigr{|}$. A classical result of Kawada and Itô [KI40] states that a nonzero boolean function $f:G\to\{0,1\}$ satisfies $\left|\!\left|f\right|\!\right|_{\rm A}\leq 1$ if and only if $f=\mathop{\mathbf{1}}\nolimits_{s+H}$ for a coset $s+H$ of $G$. The quantitative idempotent theorem states that every $f:G\to\{0,1\}$ with $\left|\!\left|f\right|\!\right|_{\rm A}\leq\gamma$ can be expressed as a signed sum
$f=\sum_{i=1}^{L}\pm\mathop{\mathbf{1}}\nolimits_{s_{i}+H_{i}}$ (1)
for some $L\leq\exp(\gamma^{3+o(1)})$ and some cosets $s_{1}+H_{1},\ldots,s_{L}+H_{L}$ of $G$. The term “idempotent” refers to the fact that boolean functions are precisely those that satisfy $f^{2}=f$, and therefore, Green and Sanders’ theorem characterizes the idempotents that have small Fourier algebra norm.
Blocky matrices and Schur multipliers.
The aim of this paper is to investigate an analogue of the program surrounding Cohen’s idempotent theorem in the setting of Schur multipliers, which we now introduce.
Let $\ell^{2}$ denote the Hilbert space of all square summable complex sequences, and let $B(\ell^{2})$ be the collection of all bounded linear operators $A:\ell^{2}\to\ell^{2}$ equipped with its operator norm
$\left|\!\left|A\right|\!\right|_{\rm op}=\sup_{\begin{subarray}{c}y\in\ell^{2}\\ y\neq 0\end{subarray}}\frac{\left|\!\left|Ay\right|\!\right|_{2}}{\left|\!\left|y\right|\!\right|_{2}}.$
Every operator $A\in B(\ell^{2})$ is uniquely identified by its associating matrix $(a_{i,j})_{i,j\in\mathbf{N}}$ defined by $a_{i,j}=\bigl{\langle}Ae_{i},e_{j}\bigr{\rangle}$, where $\{e_{i}\}_{i\in\mathbf{N}}$ is the standard orthonormal basis of $\ell^{2}$.
Every matrix $M:\mathbf{N}\times\mathbf{N}\to\mathbf{C}$ gives rise to a linear transformation on the space of all matrices $A:\mathbf{N}\times\mathbf{N}\to\mathbf{C}$ via the map $S_{M}:A\mapsto M\circ A$, where $\circ$ denotes the entrywise (Schur) product. The matrix $M$ is called a Schur multiplier if $S_{M}$ maps $B(\ell^{2})$ into itself, that is $M\circ A\in B(\ell^{2})$ for every $A\in B(\ell^{2})$. Equivalently, $M$ is a Schur multiplier if its Schur multiplier norm, defined by
$\left|\!\left|M\right|\!\right|_{\rm m}=\sup_{\begin{subarray}{c}A\in B(\ell^{2})\\ A\neq 0\end{subarray}}\frac{\left|\!\left|M\circ A\right|\!\right|_{\rm op}}{\left|\!\left|A\right|\!\right|_{\rm op}},$ (2)
is finite. Since the Schur multiplier norm satisfies
$\left|\!\left|A\circ B\right|\!\right|_{\rm m}\leq\left|\!\left|A\right|\!\right|_{\rm m}\cdot\left|\!\left|B\right|\!\right|_{\rm m},$ (3)
the set of Schur multipliers is closed under addition and Schur product, and thus forms a Banach algebra.
An element $a$ of an algebra is said to be idempotent if $a^{2}=a$. Any matrix satisfying $M\circ M=M$ must be boolean, but not all infinite boolean matrices are Schur multipliers. A boolean Schur multiplier $M:\mathbf{N}\times\mathbf{N}\to\{0,1\}$ corresponds to the natural operation that maps any matrix $A$ to a new matrix that agrees with $A$ on the support of $M$ and is zero everywhere else. The following is the central question of interest in this article:
What are the idempotent elements of the Schur multiplier algebra?
As a starting point, we discuss a simple characterization of the contractive idempotent elements, i.e., those with Schur multiplier norm at most 1.
We call a Boolean matrix $B:\mathbf{N}\times\mathbf{N}\to\{0,1\}$ blocky if there exist families $\{S_{i}\}_{i\in\mathbf{N}}$ and $\{T_{i}\}_{i\in\mathbf{N}}$ of pairwise disjoint subsets of $\mathbf{N}$ such that the support of $B$ is exactly $\bigcup_{i\in\mathbf{N}}S_{i}\times T_{i}$. Simple examples of blocky matrices are zero matrices, all-$1$s matrices, and identity matrices.
Proposition 1.1 ([Liv95]).
If a boolean matrix $A$ satisfies $\left|\!\left|A\right|\!\right|_{\rm m}\leq 1$, then $A$ is a blocky matrix. In particular, either $A=0$ and $\left|\!\left|A\right|\!\right|_{\rm m}=0$, or $A$ is a nonzero blocky matrix and $\left|\!\left|A\right|\!\right|_{\rm m}=1$.
Proof.
It is straightforward to verify that the Schur multiplier norm of every non-zero blocky matrix is $1$. To prove the converse, we must show that a boolean matrix with $\left|\!\left|A\right|\!\right|_{\rm m}\leq 1$ does not contain a $2\times 2$ submatrix with exactly three $1$-entries. A direct calculation, due to L. Livshits [Liv95], shows that (${1}{0}{1}\big{)}\right|\!\right|_{\rm m}=2/\sqrt{3}>1$ 1, and the Schur multiplier norm does not increase upon restriction to a submatrix. ∎
Consider the algebra of functions $f:G\to\mathbf{C}$ under addition and pointwise multiplication. This is a Banach algebra under the norm $\left|\!\left|\cdot\right|\!\right|_{\rm A}$, since $\left|\!\left|fg\right|\!\right|_{\rm A}\leq\left|\!\left|f\right|\!\right|_{\rm A}\cdot\left|\!\left|g\right|\!\right|_{\rm A}$. The idempotent elements of this algebra are precisely the boolean functions $f:G\to\{0,1\}$, and Proposition 1.1 is the matrix analogue of Kawada and Itô’s theorem that a function is idempotent and contractive if and only if it is the indicator of a coset.
In the same way that Cohen’s idempotent theorem shows that any idempotent element of the Fourier algebra can be written as a finite sum of contractive idempotents, it is an open problem, first stated (to our knowledge) in [KP05, Section 2], whether any idempotent Schur multiplier can be written as a finite sum of contractive idempotents in the algebra of Schur multipliers. This question is considered one of the challenging open problems in the area (see [TT10] and [ELT16, Question 3.13]). A straightforward compactness argument, presented in [HHH23, Theorem 3.10], shows that a positive resolution to this problem, while only being meaningful for infinite matrices, is equivalent to the following conjecture for finite matrices.
Conjecture 1.2 ([HHH23], Conjecture III).
Suppose that $A$ is a finite boolean matrix with $\left|\!\left|A\right|\!\right|_{\rm m}\leq\gamma$. Then we may write
$A=\sum_{i=1}^{L}\sigma_{i}B_{i},$ (4)
where the $\sigma_{i}$ are signs, the $B_{i}$ are blocky matrices, and $L$ depends only on $\gamma$.
1.2 is closely related to the quantitative version of Cohen’s idempotent theorem. Given a finite group $G$ and a function $f:G\to\mathbf{C}$, define the matrix $M_{f}:G\times G\to\mathbf{C}$ by setting $M_{f}(x,y)=f(y^{-1}x)$. We have (see, e.g., [HHH23, Corollary 3.13])
$\left|\!\left|M_{f}\right|\!\right|_{\rm m}=\left|\!\left|f\right|\!\right|_{\rm A}.$ (5)
Therefore, it easily follows from the Green–Sanders theorem (and its extension to nonabelian groups by Sanders) that 1.2 is true for convolution matrices, i.e., those of the form $M_{f}(x,y)=f(y^{-1}x)$ for some boolean function $f:G\to\{0,1\}$.
We define the block complexity $\operatorname{block}(A)$ of an $m\times n$ (integer) matrix $A$ to be the smallest integer $L$ such that there exist blocky matrices $B_{1},\ldots,B_{L}$ and signs $\sigma_{1},\ldots,\sigma_{L}$ with $A=\sum_{i=1}^{L}\sigma_{i}B_{i}$. D. Avraham and A. Yehudayoff [AY24] proved that the block complexity of a random $n\times n$ boolean matrix is at least $n/\bigl{(}4\log_{2}(2n)\bigr{)}$ with high probability.
It is immediate from the definition of the block complexity that
$\left|\!\left|A\right|\!\right|_{\rm m}\leq\operatorname{block}(A).$
1.2 claims that, conversely, the block complexity can be bounded by a function of the Schur multiplier norm. In this paper, we prove that any Boolean matrix with bounded Schur multiplier norm has block complexity at most polylogarithmic in its dimension. All logarithms in this paper are understood to be natural, unless otherwise noted.
Theorem 1.3 (Main theorem).
Let $A$ be an $m\times n$ integer matrix with $\left|\!\left|A\right|\!\right|_{\rm m}\leq\gamma$. Denoting $k=\min\{m,n\}$, the block complexity of $A$ satisfies
$\operatorname{block}(A)\leq 2^{O(\gamma^{7})}\log\bigl(k\bigr{missing})^{2}.$ (6)
Note that, similar to Cohen’s idempotent theorem, Theorem 1.3 holds more generally for integer matrices.
The factorization norm.
In the proof of Theorem 1.3, we will mainly work with Grothendieck’s characterization of Schur multipliers through the $\gamma_{2}$ factorization norm. The $\gamma_{2}$ factorization norm (or the $\gamma_{2}$ norm for short) of a complex matrix $A$ is defined as
$|\!|A|\!|_{\gamma_{2}}=\min_{UV=A}\left|\!\left|U\right|\!\right|_{\rm row}\left|\!\left|V\right|\!\right|_{\rm col},$ (7)
where the minimum is taken over all factorizations $UV$ of $A$, $\left|\!\left|U\right|\!\right|_{\rm row}$ is the largest $\ell^{2}$-norm of a row in $U$ and $\left|\!\left|V\right|\!\right|_{\rm col}$ is the largest $\ell^{2}$-norm of a column in $V$. Note that the $\gamma_{2}$ norm cannot increase by restricting to a submatrix, since removing rows of $U$ or columns of $V$ cannot increase $\left|\!\left|U\right|\!\right|_{\rm row}$ or $\left|\!\left|V\right|\!\right|_{\rm col}$. In particular, $|\!|A|\!|_{\gamma_{2}}$ is at least $|\!|A|\!|_{\rm max}=\max_{(x,y)\in X\times Y}\bigl{|}A(x,y)\bigr{|}$.
The algebra of Schur multipliers is germane to the study of the factorization norm, because it turns out that $\left|\!\left|A\right|\!\right|_{\rm m}=|\!|A|\!|_{\gamma_{2}}$ for every matrix $A$; this result is due to A. Grothendieck [Gro54] (see also [Pis12] for a proof in a more modern context).
Definitions and notation.
We shall often find it convenient to regard a matrix as a function with a set $X\times Y$ as its domain. In the rest of the paper, $X$ and $Y$ will be assumed finite unless otherwise stated, and we reserve in advance the variables $m=|X|$ and $n=|Y|$. Then, when we write $A\in\mathbf{Z}^{X\times Y}$ (for instance), we mean that $A$ is an $m\times n$ integer matrix with rows indexed by $X$ and columns indexed by $Y$. If $A$ is a matrix with domain $X\times Y$ and let $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$, we may write $A_{X^{\prime}\times Y^{\prime}}$ for the matrix on the domain $X^{\prime}\times Y^{\prime}$ with $A_{X^{\prime}\times Y^{\prime}}(x,y)=A(x,y)$ for all $x\in X^{\prime}$ and $y\in Y^{\prime}$.
Now suppose that $A$ is a complex matrix with $|\!|A|\!|_{\gamma_{2}}\leq\gamma$. By rescaling the vectors given by the definition of factorization norm, we may express $A$ as a product $A=UV$ for some matrices $U$ and $V$ such that
$\max\bigl{(}\left|\!\left|u_{1}\right|\!\right|_{2},\ldots,\left|\!\left|u_{m}\right|\!\right|_{2}\bigr{)}\leq 1$
and
$\max\bigl{(}\left|\!\left|v_{1}\right|\!\right|_{2},\ldots,\left|\!\left|v_{n}\right|\!\right|_{2}\bigr{)}\leq\gamma,$
where $u_{1},\ldots,u_{m}$ are the rows of $U$ and $v_{1},\ldots,v_{n}$ are the columns of $V$. We shall call such a factorization of $A$ a $\gamma$-factorization. If $U$ is $m\times t$ and $V$ is $t\times n$ for some integer $t$, note that the set $\{u_{1},\ldots,u_{m},v_{1},\ldots,v_{n}\}$ is contained in a $t^{\prime}$-dimensional subspace of $\mathbf{C}^{t}$ for some $t^{\prime}\leq m+n$. So, without loss of generality, we always assume that all the vectors in a $\gamma$-factorization of an $m\times n$ matrix belong to $\mathbf{C}^{m+n}$.
2 The Littlestone dimension
The Littlestone dimension is a combinatorial parameter of sign matrices that characterizes regret bounds in online learning [Lit88]. Towards our development of a generalization of the Littlestone dimension for real matrices, we devote this section to recalling some properties of the ordinary Littlestone dimension.
We begin with the definition of the Vapnik–Chervonenkis (VC) dimension. The VC dimension of a sign matrix $A\in\{-1,1\}^{X\times Y}$ is the size of the largest subset $X^{\prime}$ of $X$ with the following property: for every $b:X^{\prime}\to\{-1,1\}$, there exists a column $y$ such that $A(x,y)=b(x)$ for all $x\in X^{\prime}$. When this condition holds, we say that $A$ shatters the set $X^{\prime}$.
The Littlestone dimension relaxes this definition by shattering decision trees instead of sets. A mistake tree of depth $d$ over a domain $X$ is a complete binary tree of depth $d$ in which
  • i)
    each internal node $\nu$ is labelled with an element $x(\nu)\in X$; and
  • ii)
    each edge $e$ is labelled with a sign $\sigma(e)\in\{-1,1\}$, where $b(e)=-1$ indicates a left child and $\sigma(e)=1$ indicates a right child.
For the purposes of the Littlestone dimension, a matrix $A\in\{-1,1\}^{X\times Y}$ is said to shatter a mistake tree over $X$ if, for every root-to-leaf path $(\nu_{1},\ldots,\nu_{d+1})$, there exists a column $y\in Y$ such that $A\bigl{(}x(\nu_{i}),y\bigr{)}=\sigma(\nu_{i}\nu_{i+1})$ for all $i\in[d]$. Then the Littlestone dimension of $A$, denoted by $\operatorname{Ldim}(A)$, is the largest integer $d$ for which there exists a mistake tree of depth $d$ that is shattered by $A$. The Littlestone dimension is at least the VC dimension, since every shattered set $X^{\prime}=\{x_{1},\ldots,x_{d}\}$ gives rise to a shattered mistake tree of depth $d$ where all nodes at level $i$ are labelled with $x_{i}$.
For any sign matrix $A$ with $\operatorname{Ldim}(A)=d$ and any $20<\epsilon<1/$ 2, one can find a fixed sign vector that matches an $\epsilon^{d}$ proportion of the columns of $A$, with error at most $\epsilon$.
Proposition 2.1.
Let $A\in\{-1,1\}^{X\times Y}$ be a matrix with $\operatorname{Ldim}(A)=d$ and let $\epsilon\in(0,1/2)$. There exists a function $\sigma:X\to\{-1,1\}$ and some subset $S\subseteq Y$ with $|S|\geq\epsilon^{d}|Y|$ such that
$\Pr_{y\in S}\bigl{[}A(x,y)\neq\sigma(x)\bigr{]}\leq\epsilon$ (8)
for all $x\in X$.
Proof.
Let $A\in\{-1,1\}^{X\times Y}$ be a matrix, and let $x\in S$ be an arbitrary row of $A$. Let
$Y_{x}^{-}=\{y\in Y:A(x,y)=-1\}\qquad\hbox{and}\qquad Y_{x}^{+}=\{y\in Y:A(x,y)=1\}.$
By the definition of the Littlestone dimension, we have
$\operatorname{Ldim}(A)=1+\min\bigl{\{}\operatorname{Ldim}(A_{X\times Y_{x}^{-}}),\operatorname{Ldim}(A_{X\times Y_{x}^{+}})\bigr{\}},$ (9)
which implies that the Littlestone dimension must drop on at least one of $A_{X\times Y_{x}^{-}}$ or $A_{X\times Y_{x}^{+}}$.
If $S=Y$ and $\sigma(x)=\operatorname{sgn}\bigl{(}\sum_{y\in S}A(x,y)\bigr{)}$ does not satisfy the condition, then there is a row $x\in X$ with
$\Pr_{y\in S}\bigl{[}A(x,y)\neq\sigma(x)\bigr{]}>\epsilon,$ (10)
which implies that each of them $Y_{x}^{-}$ and $Y_{x}^{+}$ comprises at most a $(1-\epsilon)$ proportion of $Y$. By the claim of the previous paragraph, removing all the columns in one of these two sets decreases the Littlestone dimension by $1$.
The proposition now follows by induction on $d$, since when $d=0$, each row must either contain only $-1$s or only $1$s, which forces all columns to be identical. ∎
Next, we show that for sign matrices, the Littlestone dimension provides a lower bound for (the square of) the $\gamma_{2}$ norm. The following proposition essentially follows from the classic mistake-bound analysis [Ros58] of the perceptron algorithm [MP43].
Proposition 2.2.
Every $A\in\{-1,1\}^{X\times Y}$ satisfies
$|\!|A|\!|_{\gamma_{2}}\geq\sqrt{\operatorname{Ldim}(A)}.$ (11)
Proof.
Let $\gamma=|\!|A|\!|_{\gamma_{2}}$ and $d=\operatorname{Ldim}(A)$. Consider a $\gamma$-factorization $A=UV$ of $A$, so that $\left|\!\left|u_{x}\right|\!\right|\leq 1$ for all $x\in X$ and $\left|\!\left|v_{y}\right|\!\right|\leq\gamma$ for all $y\in Y$. By the definition of the Littlestone dimension, $A$ shatters a mistake tree of depth $d$ over $X$. We shall recursively construct a root-to-leaf path $(\nu_{1},\ldots,\nu_{d+1})$ that furnishes our desired bound. Let $\nu_{1}$ be the root of the tree and set $\sigma_{1}=1$. Now for all $1\leq k\leq d$, let $\nu_{k+1}$ be the child of $\nu_{k}$ with $\sigma(\nu_{k}\nu_{k+1})=\sigma_{k}$ and set
$\sigma_{k+1}=-\operatorname{sgn}\biggl{\langle}\sum_{i=1}^{k}\sigma_{i}u_{x(\nu_{i})},u_{x(\nu_{k+1})}\biggr{\rangle}.$ (12)
For each $11\leq k\leq d-$ 1, we have
$\biggl{|}\!\biggl{|}\sum_{i=1}^{k+1}\sigma_{i}u_{x(\nu_{i})}\biggr{|}\!\biggr{|}^{2}$ $=\biggl{|}\!\biggl{|}\sum_{i=1}^{k}\sigma_{i}u_{x(\nu_{i})}\biggr{|}\!\biggr{|}^{2}+\bigl{|}\!\bigl{|}u_{x(\nu_{k+1})}\bigr{|}\!\bigr{|}^{2}+2\sigma_{k+1}\biggl{\langle}\sum_{i=1}^{k}\sigma_{i}u_{x(\nu_{i})},u_{x(v_{k+1})}\biggr{\rangle}$ (13)
$\leq\bigl{|}\!\bigl{|}u_{x(\nu_{k+1})}\bigr{|}\!\bigr{|}^{2}+1,$ (14)
by our choice of $\sigma_{k+1}$, and consequently
$\biggl{|}\!\biggl{|}\sum_{i=1}^{d}\sigma_{i}u_{x(\nu_{i})}\biggr{|}\!\biggr{|}^{2}\leq d.$ (15)
By assumption, $A$ shatters the mistake tree, so there is a column $y\in Y$ with $A\bigl{(}x(\nu_{i}),y\bigr{)}=\sigma_{i}$ for all $1\leq i\leq d$. Hence,
$d=\sum_{i=1}^{d}A\bigl{(}x(\nu_{i}),y\bigr{)}\sigma_{i}=\biggl{\langle}\sum_{i=1}^{d}\sigma_{i}u_{x(\nu_{i})},v_{y}\biggr{\rangle}\leq\sqrt{d}\left|\!\left|v_{y}\right|\!\right|\leq\gamma\sqrt{d}$ (16)
by the Cauchy–Schwarz inequality. ∎
3 A weighted Littlestone dimension for real matrices
Because the inductive step in our eventual proof of Theorem 1.3 does not preserve the boolean entries of the starting matrix, we must develop a generalization of the Littlestone dimension for real matrices. This we shall do in the present section, proving analogues of Propositions 2.1 and 2.2.
Let $\alpha>0$ be a parameter. An $\alpha$-weighted mistake tree of depth $d$ over a domain $X$ is a complete binary tree of depth $d$ in which
  • i)
    each internal node $\nu$ is labelled with an element $x(\nu)\in X$; and
  • ii)
    each edge $e$ is labelled with a real number $w(e)\in\mathbf{R}$, such that if $e_{1}$ and $e_{2}$ are the two edges connecting a vertex $\nu$ to its children, then $\bigl{|}w(e_{1})-w(e_{2})\bigr{|}\geq\alpha$.
We say that $A\in\mathbf{R}^{X\times Y}$ shatters such a weighted mistake tree if for every root-to-leaf path $(\nu_{1},\ldots,\nu_{d+1})$, there exists a column $y\in Y$ such that $A\bigl{(}x(\nu_{i}),y\bigr{)}=w(\nu_{i}\nu_{i+1})$ for all $1\leq i\leq d$. Then we define the $\alpha$-weighted Littlestone dimension, denoted $\operatorname{Ldim}_{\alpha}(A)$, to be the largest $d$ such that there exists an $\alpha$-weighted mistake tree of depth $d$ shattered by $A$.
The following proposition shows that for all real matrices $A$ with small $\operatorname{Ldim}_{\alpha}(A)$, there is a function $g$ such that the entries in each row $x$ are mostly within $2\alpha$ of $g(x)$, after restricting to a large subset of columns.
Proposition 3.1.
Let $\alpha$ and $\epsilon$ be positive parameters. Suppose that $A\in\mathbf{R}^{X\times Y}$ is a matrix with $\operatorname{Ldim}_{\alpha}(A)=d$ and $|\!|A|\!|_{\rm max}=M$. There exists a subset $S\subseteq Y$ with
$|S|\geq|Y|\biggl{(}\frac{\epsilon\alpha}{2M}\biggr{)}^{d}$ (17)
and a function $g:X\to[-M,M]$ such that for every $x\in X$,
$\Pr_{y\in Y}\Bigl{[}\bigl{|}A(x,y)-g(x)\bigr{|}\geq 2\alpha\Bigr{]}\leq\epsilon.$ (18)
Proof.
If for every $x\in X$ there is some $k\in\mathbf{N}$ with
|$\biggl{|}\Bigl{\{}y\in Y:A(x,y)\in\bigl{[}-M+(k-2)\alpha,-M+(k+2)\alpha\bigr{]}\Bigr{\}}\biggr{|}\geq(1-\epsilon)|Y|,$ (19)
then we can set $g(x)=-M+k\alpha$ and $S=Y$ to prove the assertion. So, assume that there is some $x\in X$ for which such a $k$ does not exist. Letting
$S_{i}=\Bigl{\{}y\in Y:A(x,y)\in\bigl{[}-M+(i-1)\alpha,-M+i\alpha\bigr{]}\Bigr{\}},$ (20)
this assumption can be re-expressed as
$|S_{k-1}\cup S_{k}\cup S_{k+1}\cup S_{k+2}|<(1-\epsilon)|Y|$ (21)
for all $k\in\mathbf{N}$. Since $A(x,y)\in[-M,M]$ for all $y\in Y$, there must exist some $i\in\mathbf{N}$ with $|S_{i}|\geq\alpha|Y|/(2M)$. On the other hand, we know that
$|S_{i-1}\cup S_{i}\cup S_{i+1}\cup S_{i+2}|\leq(1-\epsilon)|Y|,$ (22)
so there must be some $j\in\mathbf{N}$ with $|j-i|\geq 2$ such that $|S_{j}|\geq\epsilon\alpha|Y|/(2M)$. Note that every $y\in S_{i}$ and $y^{\prime}\in S_{j}$ satisfy $\bigl{|}A(x,y)-A(x,y^{\prime})\bigr{|}\geq\alpha$, so we see that
$\operatorname{Ldim}_{\alpha}(A)\geq 1+\min\bigl{\{}\operatorname{Ldim}_{\alpha}(A_{X\times S_{i}}),\operatorname{Ldim}_{\alpha}(A_{X\times S_{j}})\bigr{\}}.$ (23)
Hence we have $\operatorname{Ldim}_{\alpha}(A_{X\times S_{i}})\leq d-1$ or $\operatorname{Ldim}_{\alpha}(A_{X\times S_{j}})\leq d-1$, and may now perform induction on $d$.
When $d=0$, pick any $y_{0}\in Y$ and let $g(x)=A(x,y_{0})$ for all $x\in X$. The fact that $A$ does not shatter any $\alpha$-weighted mistake tree of positive depth tells us that $A(x,y)\in\bigl{(}g(x)-\alpha,g(x)+\alpha)$ for all $(x,y)\in X\times Y$, and hence we may take $S=Y$. ∎
Like the ordinary Littlestone dimension, our $\alpha$-weighted Littlestone dimension is bounded polynomially in terms of the $\gamma_{2}$-norm (up to a constant factor that depends on $\alpha$).
Proposition 3.2.
Let $A\in\mathbf{R}^{X\times Y}$ be a matrix with $\operatorname{Ldim}_{\alpha}(A)=d$ (for some $\alpha>0$), $|\!|A|\!|_{\gamma_{2}}=\gamma$, and $|\!|A|\!|_{\rm max}=M$. Then
$\gamma\geq\frac{\alpha\sqrt{d}}{2(M+1)}-1,$ (24)
and consequently
$d=O_{\alpha}(\gamma^{4}).$ (25)
Proof.
Consider a $\gamma$-factorization $A=UV$, where $U$ has rows $u_{x}$ indexed by $x\in X$ and $V$ has columns $v_{y}$ indexed by $y\in Y$. Fix an $\alpha$-weighted mistake tree of depth $d$ shattered by $A$. For each internal node $\nu$ in the tree with edges $e$ and $e^{\prime}$ connecting it to its children, we define
$\widetilde{u}(v)=u_{x(\nu)}\oplus\frac{w(e)+w(e^{\prime})}{2}$ (26)
(where, for a vector $u\in\mathbf{R}^{t}$ and a real number $\beta$, $u\oplus\beta$ denotes the vector in $\mathbf{R}^{t+1}$ obtained by appending $\beta$ to $u$). Note that
$\left|\!\left|\widetilde{u}(\nu)\right|\!\right|\leq\left|\!\left|u_{x(\nu)}\right|\!\right|+|\!|A|\!|_{\rm max}\leq M+1.$
As in the proof of Proposition 2.2, we recursively define a root-to-leaf path that supplies our final bound. We let $\nu_{1}$ denote the root of the tree and set $\sigma_{1}=1$. For $1\leq k\leq d$, let $\nu$ and $\nu^{\prime}$ be the two children of $\nu_{k}$, where we may assume that $w(\nu_{k}\nu)\geq\alpha+w(\nu_{k}\nu^{\prime})$ without losing generality. Let
$\sigma_{k+1}=-\operatorname{sgn}\biggl{\langle}\sum_{i=1}^{k}\sigma_{i}\widetilde{u}(\nu_{i}),\widetilde{u}(\nu_{k+1})\biggr{\rangle}.$ (27)
If $\sigma_{k+1}=1$, set
$\nu_{i+1}=v\qquad\hbox{and}\qquad\nu_{i+1}^{\prime}=\nu^{\prime}$
and if $\sigma_{k+1}=-1$, set
$\nu_{i+1}=\nu^{\prime}\qquad\hbox{and}\qquad\nu_{i+1}^{\prime}=\nu.$
Either way, we have
$\sigma_{k+1}\bigl{(}w(\nu_{i}\nu_{i+1})-w(\nu_{i}\nu_{i+1}^{\prime})\bigr{)}\geq\alpha.$ (28)
By (27), for all $11\leq k\leq d-$ 1,
$\biggl{|}\!\biggl{|}\sum_{i=1}^{k+1}\sigma_{i}\widetilde{u}(\nu_{i})\biggr{|}\!\biggr{|}^{2}$ $=\biggl{|}\!\biggl{|}\sum_{i=1}^{k}\sigma_{i}\widetilde{u}(\nu_{i})\biggr{|}\!\biggr{|}^{2}+\bigl{|}\!\bigl{|}\widetilde{u}(\nu_{i})\bigr{|}\!\bigr{|}^{2}+2\sigma_{k+1}\biggl{\langle}\sum_{i=1}^{k}\sigma_{i}\widetilde{u}(\nu_{i}),\widetilde{u}(\nu_{k+1})\biggr{\rangle}$
$\leq\biggl{|}\!\biggl{|}\sum_{i=1}^{k}\sigma_{i}\widetilde{u}(\nu_{i})\biggr{|}\!\biggr{|}^{2}+\bigl{|}\!\bigl{|}\widetilde{u}(\nu_{k+1})\bigr{|}\!\bigr{|}^{2},$
which shows that
$\biggl{|}\!\biggl{|}\sum_{i=1}^{k}\sigma_{i}\widetilde{u}(\nu_{i})\biggr{|}\!\biggr{|}^{2}\leq\sum_{i=1}^{d}\bigl{|}\!\bigl{|}\widetilde{u}(\nu_{i})\bigr{|}\!\bigr{|}^{2}\leq d(M+1)^{2}.$ (29)
On the other hand, since $A$ shatters the tree, there exists a column $y\in Y$ with
$\bigl{\langle}u_{x(\nu_{i})},v_{y}\bigr{\rangle}=A\bigl{(}x(\nu_{i}),y\bigr{)}=w(\nu_{i}\nu_{i+1})$ (30)
for all $1\leq i\leq d$. Let $\widetilde{v}=v_{y}\oplus-1$. Now by (28), we bound
$\biggl{\langle}\sum_{i=1}^{d}\sigma_{i}u_{i},\widetilde{v}\biggr{\rangle}$ $=\sum_{i=1}^{d}\sigma_{i}\biggl{(}\bigl{\langle}u_{x(\nu_{i})},v_{y}\bigr{\rangle}-\frac{w(\nu_{i}\nu_{i+1})+w(\nu_{i}\nu_{i+1}^{\prime})}{2}\biggr{)}$ (31)
$=\sum_{i=1}^{d}\sigma_{i}\biggl{(}\frac{w(\nu_{i}\nu_{i+1})-w(\nu_{i}\nu_{i+1}^{\prime})}{2}\biggr{)}$ (32)
$\geq\frac{\alpha d}{2}.$ (33)
By the Cauchy–Schwarz inequality, we see that
$\frac{\alpha d}{2}\leq\biggl{|}\!\biggl{|}\sum_{i=1}^{d}b_{i}\widetilde{u}(\nu_{i})\biggr{|}\!\biggr{|}\cdot\bigl{|}\!\bigl{|}\widetilde{v}\bigr{|}\!\bigr{|}\leq\sqrt{d}(M+1)(1+\gamma).$ (35)
Hence
$\gamma\geq\frac{\alpha\sqrt{d}}{2(M+1)}-1,$ (36)
and we also have $d=O(\gamma^{4}/\alpha^{2})$, since $|\!|A|\!|_{\rm max}\leq|\!|A|\!|_{\gamma_{2}}$. ∎
4 Three lemmas
In this section, we establish three miscellaneous lemmas. We begin with a simple upper bound on the block complexity of an integer matrix in terms of the $\ell_{1}$ norms of its rows.
Lemma 4.1.
If $A\in\mathbf{Z}^{X\times Y}$ contains only nonnegative entries, then
$\operatorname{block}(A)\leq\max_{x\in X}\sum_{y\in Y}\bigl{|}A(x,y)\bigr{|}.$ (37)
More generally, every $A\in\mathbf{Z}^{X\times Y}$ satisfies
$\operatorname{block}(A)\leq 2\max_{x\in X}\sum_{y\in Y}\bigl{|}A(x,y)\bigr{|}.$ (38)
Proof.
The second statement follows from the first by splitting a matrix $A$ into $A_{+}-A_{-}$, where both $A_{+}$ and $A_{-}$ are nonnegative, and then observing that $\operatorname{block}(A)\leq\operatorname{block}(A_{+})+\operatorname{block}(A_{-})$.
Hence, we may assume that all entries of $A$ are nonnegative. Let $X^{\prime}$ be the set of nonzero rows of $A$. For every nonzero row $x\in X^{\prime}$, choose $y(x)\in Y$ such that $A\bigl{(}x,y(x)\bigr{)}\neq 0$. The set $\bigl{\{}\bigl{(}x,y(x)\bigr{)}:x\in X^{\prime}\bigr{\}}$ defines a blocky matrix $B\in\{0,1\}^{X\times Y}$. Furthermore, all the entries in $A^{\prime}=A-B$ are nonnegative, and
$\max_{x\in X}\sum_{y\in Y}A^{\prime}(x,y)=\max_{x\in X}\sum_{y\in Y}A(x,y)-1.$ (39)
This argument may be iterated until $A$ is the zero matrix. ∎
Next, we prove a simple technical lemma, showing that if some set of bounded vectors has a large average vector $\widehat{v}$, then subtracting $\widehat{v}$ from many of the individual vectors significantly reduces their norms.
Lemma 4.2.
Let $v_{1},\ldots,v_{r}$ be vectors in a Hilbert space, and let $\widehat{v}=\mathop{\mathbf{E}}\nolimits_{i\in[r]}v_{i}$ denote their average. If $\left|\!\left|v_{i}\right|\!\right|\leq\gamma$ for all $1\leq i\leq r$, and $\left|\!\left|\widehat{v}\right|\!\right|=c$, there exists a subset $S\subseteq[r]$ with $|S|\geq c^{2}r/(2\gamma^{2})$ such that for every $i\in S$,
$\left|\!\left|v_{i}-\widehat{v}\right|\!\right|^{2}\leq\left|\!\left|v_{i}\right|\!\right|^{2}-\frac{c^{2}}{2}.$ (40)
Proof.
Define
$S=\bigl{\{}i\in[r]:\left|\!\left|v_{i}-\widehat{v}\right|\!\right|^{2}\leq\left|\!\left|v_{i}\right|\!\right|^{2}-c^{2}/2\bigr{\}},$ (41)
and observe that
$\mathop{\mathbf{E}}\nolimits_{i\in[r]}\left|\!\left|v_{i}-\widehat{v}\right|\!\right|^{2}=\mathop{\mathbf{E}}\nolimits_{i\in[r]}\left|\!\left|v_{i}\right|\!\right|^{2}-\left|\!\left|\widehat{v}\right|\!\right|^{2}=\mathop{\mathbf{E}}\nolimits_{i\in[r]}\left|\!\left|v_{i}\right|\!\right|^{2}-c^{2}.$ (42)
Therefore,
$c^{2}=\mathop{\mathbf{E}}\nolimits_{i\in[r]}\left|\!\left|v_{i}\right|\!\right|^{2}-\left|\!\left|v_{i}-\widehat{v}\right|\!\right|^{2}\leq\frac{c^{2}}{2}+\frac{|S|}{r}\gamma^{2}.$ (43)
This gives
$\frac{|S|}{r}\geq\frac{c^{2}}{2\gamma^{2}},$ (44)
which is what we wanted. ∎
For a boolean matrix $A\in\{0,1\}^{X\times Y}$ with $|Y|=n$, the third and final lemma of this section supplies a partition $Y=\bigcup_{i=1}^{k}S_{i}$ such that
  • i)
    for every $S_{i}$ there is a row $x_{i}$ such that $A(x_{i},y)=1$ for all $y\in S_{i}$; and
  • ii)
    for every row $x$ and every $\delta>0$, there are at most $O_{\delta}(\log n)$ sets $S_{i}$ with
    $\Pr_{y\in S_{i}}\bigl{[}A(x,y)=1\bigr{]}\geq\delta.$ (45)
In fact, we may establish a more general version of this statement for integer-valued matrices.
Lemma 4.3.
Let $A\in\mathbf{Z}^{X\times Y}$ be an integer matrix with no all-zero columns. There exists a partition $Y=\bigcup_{i=1}^{k}S_{i}$, along with rows $x_{1},\ldots,x_{k}\in X$ and nonzero integers $b_{1}\ldots,b_{k}\in\mathbf{Z}\setminus\{0\}$ such that
  • i)
    for every $1\leq i\leq k$, we have $A(x_{i},y)=b_{i}$ for all $y\in S_{i}$; and
  • ii)
    for all $\delta>0$, every $x\in X$, and every $b\in\mathbf{Z}\setminus\{0\}$, the number of indices $1\leq i\leq k$ for which
    $\Pr_{y\in S_{i}}\bigl{[}A(x,y)=b\bigr{]}\geq\delta$ (46)
    is at most $(\log|Y|+1)/\delta$.
Proof.
Consider the following greedy algorithm.
Algorithm I (Construct sets $S_{i}$). Given an integer matrix $A\in\mathbf{Z}^{X\times Y}$, this algorithm outputs a partition $Y=\bigcup_{i=1}^{k}S_{i}$, as well as rows $x_{1},\ldots,x_{k}$ and nonzero integers $b_{1},\ldots,b_{k}$.
  • I1.
    [Initialize.] Set $R\leftarrow Y$ and $i\leftarrow 1$.
  • I2.
    [Loop.] While $R\neq\emptyset$, repeat steps I3 and I4.
    • I3.
      [Greedy choice.] Choose $(x_{i},b_{i})\in X\times\bigl{(}\mathbf{Z}\setminus\{0\}\bigr{)}$ that maximizes the size of
      $S_{i}=\bigl{\{}y\in R:A(x_{i},y)=b_{i}\bigr{\}}.$
    • I4.
      [Update.] Set $R\leftarrow R\setminus S_{i}$ and increment $i\leftarrow i+1$.
The claim is that the output of Algorithm I satisfies our two desired properties. It is clear that the sets $S_{i}$ form a partition of $Y$, and by their definition in step I3, they satisfy property (i).
To prove the second assertion, we assign the cost $c(y)=1/|S_{i}|$ to each $y\in S_{i}$. Let $x\in X$ and $b\in\mathbf{Z}\setminus\{0\}$ be arbitrary, and consider the set
$\{y_{1},\ldots,y_{t}\}=\bigl{\{}y\in Y:A(x,y)=b\bigr{\}},$
which we have ordered such that $c(y_{1})\leq\cdots\leq c(y_{t})$.
First we show that for every $1\leq i\leq t$,
$c(y_{i})\leq\frac{1}{t-i+1}.$ (47)
Indeed, suppose towards a contraction that $c(y_{i})>1/(t-i+1)$. Let $S_{r}$ be the member of our partition with $y_{i}\in S_{r}$, and suppose that $|S_{r}|=k$, so that $1/k=c(y_{i})>1/(t-i+1)$. The set $S^{\prime}=\{y_{j}\in Y:j\geq i\}$ has size $t-i+1>k$, and $A(x,y)=b$ for all $y\in S^{\prime}$, hence by the greedy choice made in step I3, the partition element $S_{r}$ containing $y_{i}$ must be at least as large as $S^{\prime}$, and contain strictly more than $k$ elements. This is a contradiction.
Hence we bound
$\sum_{i=1}^{k}\Pr_{y\in S_{i}}\bigl{[}A(x,y)=b\bigr{]}=\sum_{i=1}^{k}\frac{\bigl{\{}y\in S_{i}:A(x,y)=b\bigr{\}}}{|S_{i}|}=\sum_{i=1}^{t}c(y_{i})\leq\sum_{i=1}^{t}\frac{1}{i}\leq\log|Y|+1,$ (48)
and the number of indices $i$ with $\Pr_{y\in S_{i}}\bigl{[}A(x,y)=b\bigr{]}\geq\delta$ can be at most $(\log|Y|+1)/\delta$. ∎
5 The main argument
We are now ready to prove our main theorem. Given a matrix $A\in\mathbf{Z}^{X\times Y}$ with $|\!|A|\!|_{\gamma_{2}}\leq\gamma$, we wish to express it as a sum of matrices with small block complexity, which in turn gives a bound on the block complexity of $A$. Our strategy shall be to partition the columns $Y$ into subsets $S_{i}$ using Lemma 4.3 and then define a new matrix $\widetilde{A}$ by averaging the columns $y$ belonging to each $S_{i}$. The block complexity of $\widetilde{A}$ can be bounded by invoking Lemma 4.1, and Lemma 4.2 will help us show that $|\!|A-\widetilde{A}|\!|_{\gamma_{2}}^{2}$ decreases by a constant, making the argument suitable for iteration.
The catch is that subtracting averages takes us outside the realm of integer-valued matrices, so we must consider a slightly wider class of matrices. This motivates the following definition, which is our analogue of the almost integer-valued functions used in [GS08a]. Given a matrix $A\in\mathbf{R}^{X\times Y}$, let $A_{\mathbf{Z}}\in\mathbf{Z}^{X\times Y}$ denote the entrywise rounding of $A$ to the nearest integer matrix, where $b+1/2$ is rounded down to $b$ for all $b\in\mathbf{Z}$. For any real $\epsilon>0$, we say that a real-valued matrix $A\in\mathbf{R}^{X\times Y}$ is $\epsilon$-almost integer-valued if $|\!|A-A_{\mathbf{Z}}|\!|_{\rm max}\leq\epsilon$.
The following lemma is the heart of our induction, allowing us to decrease the factorization norm.
Lemma 5.1 (Key lemma).
Let $A\in\mathbf{R}^{X\times Y}$ be a real-valued matrix with $|\!|A|\!|_{\gamma_{2}}=\gamma$. Suppose further that $A$ is $\epsilon$-almost integer-valued for $\epsilon=2^{-20\gamma^{2}}$. If $A_{\mathbf{Z}}$ is not an all-zero matrix, then there exists a $2\epsilon$-almost integer-valued matrix $A^{\prime}\in\mathbf{R}^{X\times Y}$ such that
$|\!|A-A^{\prime}|\!|_{\gamma_{2}}^{2}\leq\gamma^{2}-\frac{1}{8}$ (49)
and
$\operatorname{block}(A^{\prime}_{\mathbf{Z}})\leq 2^{O(\gamma^{7})}\bigl{(}\log|Y|\bigr{)}^{2}.$ (50)
Proof.
Let $M=|\!|A|\!|_{\rm max}$. If $A_{\mathbf{Z}}$ is not an all-zero matrix, then $M\geq 1/2$ and we have $\gamma\geq 1/2$ as well. This implies that $\epsilon\leq 2^{-10\gamma}\leq 1/32$. We fix a $\gamma$-factorization $A=UV$ of $A$, with $\left|\!\left|u_{x}\right|\!\right|\leq 1$ for all $x\in X$ and $\left|\!\left|v_{y}\right|\!\right|\leq\gamma$ for all $y\in Y$.
By applying Lemma 4.3 to (the nonzero columns of) $A_{\mathbf{Z}}$, we obtain rows $x_{1},\ldots,x_{k}\in X$, integers $b_{1},\ldots,b_{k}\in\mathbf{Z}\setminus\{0\}$, and sets $S_{1},\ldots,S_{k}\subseteq Y$ such that for every $1\leq i\leq k$ and every $y\in S_{i}$, we have $A_{\mathbf{Z}}(x_{i},y)=b_{i}$. (The second statement of Lemma 4.3 will not be used until later, and we defer our choice of $\delta$ until then.)
We now invoke the properties of the $\alpha$-weighted Littlestone dimension. By Proposition 3.2, we have $\operatorname{Ldim}_{1/8}(A)\leq O(\gamma^{4})$, and as a result, we also have $\operatorname{Ldim}_{1/8}(A_{X\times S_{i}})=O(\gamma^{4})$ for all $1\leq i\leq k$. Set $\epsilon_{1}=\epsilon/(10\gamma)$ and apply Proposition 3.1 to each matrix $A_{X\times S_{i}}$ with this parameter to obtain subsets $S_{i}^{\prime}$ with
$|S_{i}^{\prime}|\geq\Bigl{(}\frac{\epsilon_{1}}{\gamma}\Bigr{)}^{\!O(\gamma^{4})}|S_{i}|\geq{\epsilon_{1}}^{O(\gamma^{5})}|S_{i}|$ (51)
and functions $g^{\prime}_{i}:X\to[-M,M]$. Rounding each of these functions $g_{i}^{\prime}$ to their nearest integer functions, we obtain functions $g_{i}:X\to[-M-1/2,M+1/2]\cap\mathbf{Z}$ such that
$\Pr_{y\in S_{i}^{\prime}}\bigl{[}A_{\mathbf{Z}}(x,y)\neq g_{i}(x)\bigr{]}\leq\Pr_{y\in S_{i}^{\prime}}\Bigl{[}\bigl{|}A(x,y)-g_{i}^{\prime}(x)\bigr{|}\geq 1/4\Bigr{]}\leq\epsilon_{1}$ (52)
for all $1\leq i\leq k$, where we have applied the fact that $A$ is $\epsilon$-almost integer-valued for some $320<\epsilon\leq 1/$ 32.
For each $i$, let $\widehat{v}_{i}$ denote the average $\mathop{\mathbf{E}}\nolimits_{y\in S_{i}^{\prime}}v_{y}$, and observe that since $A_{\mathbf{Z}}(x_{i},y)=b_{i}\neq 0$,
$\bigl{|}\langle u_{x_{i}},\widehat{v}_{i}\rangle\bigr{|}=\Bigl{|}\mathop{\mathbf{E}}\nolimits_{y\in S_{i}^{\prime}}\langle u_{x_{i}},v_{y}\rangle\Bigr{|}=\Bigl{|}\mathop{\mathbf{E}}\nolimits_{y\in S_{i}^{\prime}}A(x_{i},y)\bigr{|}\geq|b_{i}-\epsilon|\geq\frac{1}{2},$ (53)
which in turn gives the bound $\left|\!\left|\widehat{v}_{i}\right|\!\right|\geq 1/2$ by the Cauchy–Schwarz inequality. We may then apply Lemma 4.2 to obtain subsets $S_{i}^{\prime\prime}\subseteq S_{i}^{\prime}$ with
$|S_{i}^{\prime\prime}|\geq\frac{1}{8\gamma^{2}}|S_{i}^{\prime}|$ (54)
such that
$\left|\!\left|v_{y}-\widehat{v}_{i}\right|\!\right|^{2}\leq\left|\!\left|v_{y}\right|\!\right|^{2}-\frac{1}{8}\leq\gamma^{2}-\frac{1}{8}$ (55)
for every $y\in S_{i}^{\prime\prime}$. Let $Y^{\prime}=\bigcup_{i=1}^{k}S_{i}^{\prime}$ and $\widetilde{Y}=\sum_{i=1}^{k}S_{i}^{\prime\prime}$; we have
$|\widetilde{Y}|\geq\frac{1}{8\gamma^{2}}|Y^{\prime}|\geq{\epsilon_{1}}^{O(\gamma^{5})}|Y|\geq 2^{-O(\gamma^{7})}|Y|.$ (56)
We now define a matrix $\widetilde{A}\in\mathbf{R}^{X\times\widetilde{Y}}$ by setting $\widetilde{v}_{y}=v_{y}-\widehat{v}_{i}$, so that we have $\left|\!\left|\widetilde{v}_{y}\right|\!\right|^{2}\leq\gamma^{2}-1/8$, and
$\widetilde{A}(x,y)=\mathop{\mathbf{E}}\nolimits_{y^{\prime}\in S_{i}^{\prime}}A(x,y^{\prime})=\langle u_{x},\widehat{v}_{i}\rangle$ (57)
for every $x\in X$ and $y\in S_{i}^{\prime\prime}$. (Notice that the average is taken over $S_{i}^{\prime}$ and not $S_{i}^{\prime\prime}$.) Then for every pair $(x,y)\in X\times S_{i}^{\prime\prime}$, we have
$A(x,y)-\widetilde{A}(x,y)=\langle u_{x},v_{y}-\widehat{v}_{i}\rangle=\langle u_{x},\widetilde{v}_{y}\rangle.$ (58)
Next, we show that $\widetilde{A}$ is $2\epsilon$-almost integer-valued and we bound the block complexity of $\widetilde{A}_{\mathbf{Z}}$. For every $x\in X$, the fact that $A$ is $\epsilon$-almost integer-valued, combined with our earlier bound $\Pr_{y\in S_{i}^{\prime}}\bigl{[}A_{\mathbf{Z}}(x,y)\neq g_{i}(x)\bigr{]}\leq\epsilon_{1}$, shows that
$\bigl{|}\widetilde{A}(x,y)-g_{i}(x)\bigr{|}=\Bigl{|}\mathop{\mathbf{E}}\nolimits_{y^{\prime}\in S_{i}^{\prime}}A(x,y^{\prime})-g_{i}(x)\Bigr{|}\leq(1-\epsilon_{1})\epsilon+\epsilon_{1}(2M+1)=\epsilon+(2\gamma+1)\epsilon_{1}\leq 2\epsilon;$ (59)
that is, $\widetilde{A}$ is a $2\epsilon$-almost integer-valued matrix, and $\widetilde{A}_{\mathbf{Z}}(x,y)=g_{i}(x)$ for all $(x,y)\in X\times S_{i}^{\prime\prime}$. Define the matrix $G\in\mathbf{Z}^{X\times[k]}$ by $G(x,i)=g_{i}(x)$. Since $\widetilde{A}_{\mathbf{Z}}$ is obtained from $G$ by creating $|S_{i}^{\prime\prime}|$ copies of each column $i$, we have $\operatorname{block}(\widetilde{A}_{\mathbf{Z}})=\operatorname{block}(G)$. Consider a fixed $x\in X$ and integer $b\neq 0$. If $g_{i}(x)=b$, then since $\Pr_{y\in S_{i}^{\prime}}\bigl{[}A_{\mathbf{Z}}(x,y)=b\bigr{]}\geq 1-\epsilon_{1}$, we have
$\Pr_{y\in S_{i}}\bigl{[}A_{\mathbf{Z}}(x,y)=b\bigr{]}\geq(1-\epsilon_{1})\frac{|S_{i}^{\prime}|}{|S_{i}|}\geq(1-\epsilon_{1}){\epsilon_{1}}^{O(\gamma^{5})}\geq 2^{-O(\gamma^{7})}.$ (60)
Now is the time to apply the second assertion of Lemma 4.3, with $\delta=2^{-O(\gamma^{7})}$. It tells us that there are at most $\bigl{(}\log|Y|+1\bigr{)}/\delta=2^{O(\gamma^{7})}\log|Y|$ such indices $i$ with $\Pr_{y\in S_{i}}\bigl{[}A_{\mathbf{Z}}(x,y)=b\bigr{]}$. In other words, for any $x\in X$, each of the nonzero integers in the range $[-M,M]$ appears in the sum $\sum_{i=1}^{k}\bigl{|}g_{i}(x)\bigr{|}$ no more than $2^{O(\gamma^{7})}\log|Y|$ times. There are no more than $2M$ nonzero integers in this range, and we have
$\sum_{i=1}^{k}\bigl{|}g_{i}(x)\bigr{|}\leq 2M^{2}\cdot 2^{O(\gamma^{7})}\log|Y|=2^{O(\gamma^{7})}\log|Y|.$ (61)
By Lemma 4.1, we see that
$\operatorname{block}(\widetilde{A}_{\mathbf{Z}})=\operatorname{block}(G)\leq 2\max_{x\in X}\sum_{i=1}^{k}\bigl{|}g_{i}(x)\bigr{|}\leq 2^{O(\gamma^{7})}\log|Y|.$ (62)
The key observation at this juncture is that, moving from a $\gamma$-factorization $A=UV$ to a $\gamma$-factorization $\widetilde{A}=\widetilde{U}\widetilde{V}$, we may take $\widetilde{U}=U$; in other words, the row vectors $u_{x}$ remain unchanged. This allows us to repeatedly apply what we derived above to the remaining columns and concatenate all the matrices $\widetilde{A}$ into a single matrix $A^{\prime}\in\mathbf{R}^{X\times Y}$ with $|\!|A-A^{\prime}|\!|_{\gamma_{2}}^{2}\leq\gamma^{2}-1/8$.
In this spirit, starting with $Y_{0}=Y$ and $\widetilde{Y}_{0}=\emptyset$, we construct a partition $Y=\bigcup_{i=1}^{t}\widetilde{Y}_{i}$ and matrices $\widetilde{A}_{i}\in\mathbf{R}^{X\times\widetilde{Y}_{i}}$ by repeatedly applying the argument above to the decreasing sequence $Y_{i}=Y\setminus\bigcup_{j=0}^{i-1}\widetilde{Y}_{j}$. Since $|Y_{i}|\leq\bigl{(}1-2^{-O(\gamma^{7})}\bigr{)}|Y_{i-1}|$, this process stops after $t\leq 2^{O(\gamma^{7})}\log|Y|$ steps.
Let $A^{\prime}\in\mathbf{R}^{X\times Y}$ be the matrix with $A^{\prime}_{X\times\widetilde{Y}_{i}}=\widetilde{A}_{i}$ for all $1\leq i\leq t$. We have
$\operatorname{block}(A^{\prime}_{\mathbf{Z}})\leq\sum_{i=1}^{t}\operatorname{block}\bigl{(}(\widetilde{A}_{i})_{\mathbf{Z}}\bigr{)}\leq t\cdot 2^{O(\gamma^{7})}\log|Y|=2^{O(\gamma^{7})}\bigl{(}\log|Y|\bigr{)}^{2}.$ (63)
By construction, the difference $A-A^{\prime}$ may be factorized $A-A^{\prime}=UV^{\prime}$ where every column $v^{\prime}_{y}$ in $V^{\prime}$ satisfies $\left|\!\left|v^{\prime}_{y}\right|\!\right|^{2}\leq\gamma^{2}-1/8$, so $|\!|A-A^{\prime}|\!|_{\gamma_{2}}^{2}\leq\gamma^{2}-1/8$. ∎
Our main theorem now follows by iterating Lemma 5.1. In fact, we prove the following more general statement for almost integer-valued matrices.
Theorem 5.2 (Generalized main theorem).
For every finite $\epsilon$-almost integer-valued matrix $A\in\mathbf{R}^{X\times Y}$ with $|\!|A|\!|_{\gamma_{2}}\leq\gamma$ and $\epsilon=2^{-20\gamma^{2}}$, the block complexity of $A_{\mathbf{Z}}$ satisfies
$\operatorname{block}(A_{\mathbf{Z}})\leq 2^{O(\gamma^{7})}\bigl{(}\log|Y|\bigr{)}^{2}.$ (64)
Proof.
The proof is by induction on $\gamma$. If $\gamma<1/2$, then as observed earlier, we have $|\!|A|\!|_{\rm max}<1/2$ and $A_{\mathbf{Z}}$ is an all-zero matrix, which has block complexity zero. Otherwise, by Lemma 5.1, there is a $2\epsilon$-almost integer-valued matrix $A^{\prime}\in\mathbf{R}^{X\times Y}$ with
$|\!|A-A^{\prime}|\!|_{\gamma_{2}}^{2}\leq\gamma^{2}-\frac{1}{8}$ (65)
and
$\operatorname{block}(A^{\prime}_{\mathbf{Z}})\leq 2^{C^{\prime}\gamma^{7}}\bigl{(}\log|Y|\bigr{)}^{2}$ (66)
for some absolute constant $C^{\prime}$. Set $C=\max\{100C^{\prime},8^{7/2}\}$. The matrix $A-A^{\prime}$ is $3\epsilon$-almost integer-valued and
$3\epsilon\leq 3\cdot 2^{-20\gamma^{2}}\leq 2^{-20(\gamma^{2}-1/8)},$ (67)
so the induction hypothesis provides the bound
$\operatorname{block}\bigl{(}(A-A^{\prime})_{\mathbf{Z}}\bigr{)}\leq 2^{C(\gamma^{2}-1/8)^{7/2}}\bigl{(}\log|Y|\bigr{)}^{2}.$ (68)
This yields
$\operatorname{block}(A_{\mathbf{Z}})$ $\leq\operatorname{block}(A^{\prime}_{\mathbf{Z}})+\operatorname{block}\bigl{(}(A-A^{\prime})_{\mathbf{Z}}\bigr{)}$ (69)
$\leq\bigl{(}2^{C^{\prime}\gamma^{7}}+2^{C(\gamma^{2}-1/8)^{7/2}}\bigr{)}\bigl{(}\log|Y|\bigr{)}^{2}$ (70)
$\leq\bigl{(}2^{C(\gamma^{2}-1/8)^{7/2}+1}\bigr{)}\bigl{(}\log|Y|\bigr{)}^{2}$ (71)
$\leq 2^{C\gamma^{7}}\bigl{(}\log|Y|\bigr{)}^{2},$ (72)
where our assumption that $\gamma\geq 1/2$ implies that $C\geq C^{\prime}\gamma^{7}/(\gamma^{2}-1/8)^{7/2}$. ∎
Concluding remarks.
There are two places in which we lose a logarithmic factor, resulting in the factor of $\bigl{(}\log|Y|\bigr{)}^{2}$ in our final theorem. The first is in Lemma 4.3, when we bound the number of indices by $\bigl{(}\log|Y|+1\bigr{)}/\delta$, and the second is in (56), where one discards a constant fraction of the columns. Tightening these bounds (especially the second) appears out of reach of our current methods; we suspect that novel ideas will be necessary to prove 1.2.
On the other hand, our polylogarithmic bound is already strong, requiring the full power of our hypothesis on $\epsilon$. Indeed, M. Göös, N. Harms, and A Riazanov recently proved [GHR25] that for every $\epsilon>0$, there exists an $n\times n$ boolean matrix $M$ and a real $\epsilon$-approximation $\widetilde{M}$ of $M$ such that $|\!|\widetilde{M}|\!|_{\gamma_{2}}\leq O_{\epsilon}(1)$, while $\left|\!\left|M\right|\!\right|_{\gamma_{2}}\geq 2^{\Omega_{\epsilon}\bigl{(}\sqrt{\log n}\bigr{)}}$ and hence $\operatorname{block}(M)\geq 2^{\Omega_{\epsilon}\bigl{(}\sqrt{\log n}\bigr{)}}$.
Related work.
A recent paper of I. Balla, L. Hambardzumyan, and I. Tomon approaches 1.2 from a different angle. To state their result, we introduce the following definition. A monochromatic rectangle in an integer matrix $A$ is the product of a subset $S$ of rows and a subset $T$ of columns such that the entries $A(i,j)$ are all the same for $(i,j)\in S\times T$. We shall call a monochromatic rectangle all of whose entries are $b$ a $b$-rectangle.
Theorem 5.3 ([BHT25], Theorem 1).
Suppose that $A$ is an $m\times n$ boolean matrix with $|\!|A|\!|_{\gamma_{2}}\leq\gamma$. There is a monochromatic rectangle $S\times T$ in $A$, where $S\subseteq[m]$ and $T\subseteq[n]$ satisfy
$\frac{|S|\cdot|T|}{mn}\geq 2^{-O(\gamma^{3})}.$ (74)
Specifically, if more than half of $A$’s entries are $1$, then $S\times T$ is a $1$-rectangle, and otherwise it is a 00-rectangle.
A proof that Theorem 5.3 would follow from 1.2 is presented as [HHH23, Lemma 3.5]. A subsequent paper of the authors uses Theorem 5.3 to show that for any matrix $A$ with bounded $\gamma_{2}$ norm, one can find a large blocky matrix that contains a constant fraction of the $1$-entries in $A$.
Theorem 5.4 ([GH25], Theorem 1.4).
Let $A$ be an $m\times n$ boolean matrix $A$ with $|\!|A|\!|_{\gamma_{2}}\leq\gamma$ in which the number of $1$-entries is $F$. There exists an $m\times n$ blocky matrix $B$ containing at least $F/2^{2^{O(\gamma)}}$ $1$-entries, such that for all $(i,j)\in[m]\times[n]$, we have $B(i,j)=1$ only if $A(i,j)=1$.
Acknowledgements
[…]